# Large Language Model for Causal Decision Making

### 29 Dec 2023 v2

@@ 2.1 Large Language Model with Tool Usage
2.2 Enhancing Pre-trained Large Language Models
Pre-trained large language models often contain billions of parameters that impose challenges when fine-tuning the entire model to adapt to downstream tasks. Two classes of research are developed to reduce such a computational burden. A straightforward but effective workaround is freezing most parts of the model and only tuning a small portion of the parameters, which includes additive methods and reparametrization-based methods (Lialin et al., 2023). Additive methods, such as adapters (Houlsby et al., 2019), augment the pre-trained model with extra parameters and only train the newly added parameters. Reparametrization-based methods, such as LoRA (Hu et al., 2021), reduce the number of trainable parameters utilizing low-rank representations. To improve the training efficiency, our proposed method utilizes the LOw-Rank Adaptation method (LoRA) (Hu et al., 2021), which reduces trainable parameters by freezing the pre-trained model and optimizing the rank-decomposition matrices.
Another line of research has demonstrated that template-based prompting can also notably enhance the performance of LLMs. Commonly used prompt types include cloze question style prompts, which consist of string templates with placeholders for the LLM to fill in (Rajagopal et al., 2021; Cui et al., 2021), and prefix style prompts that prepend tokens to the input texts of LLMs (Li and Liang, 2021; Lester et al., 2021). The latter is particularly effective for generative tasks, such as generating interpretations of numeric outputs from causal models in our LLM4Causal framework (Liu et al., 2023). A common way to create such prompts is to design them manually and such prompts usually include a few human-generated exemplars of input-output pairs for LLMs (Chung et al., 2022). These human-crafted prompts have improved performance across various tasks such as question answering and text classification (Brown et al., 2020b; Schick and Schütze, 2020). Recently, Wei et al. (2022) demonstrated that prompting language models to add reasoning steps before the actual output significantly enhances the LLM’s performance on arithmetic, commonsense, and symbolic reasoning tasks. To reduce manual effort in this process, some studies have also explored automation in prompt design. For instance, Shin et al. (2020) discretely optimized the prompt by adding tokens from a certain word collection, while Haviv et al. (2021) focuses on automatic paraphrasing of given manually designed prompts and selecting the most effective one from these variations. In our LLM4Causal framework, we utilize prefix style prompts with manually designed input-output examples to effectively collect data for fine-tuning LLMs to classify user queries into different causal tasks and interpret the numerical results from causal tools. The satisfying performance of our framework has demonstrated the effectiveness of our carefully crafted prompts.
2.3 Causal Decision Making
2.4 Large Language Models with Causal Capabilities
3 Problem Formulation
In this paper, we aim to augment existing pre-trained LLMs with the proficiency to address causal inquiries, thereby opening the door to causal decision-making processes for general audiences. As introduced previously, causal decision-making tasks can be classified into three categories: 1) causal structure learning (CSL), 2) causal effect learning (CEL), and 3) causal policy learning (CPL). Each of these primary categories consists of multiple tasks, and our study focuses on five key tasks that are particularly relevant to the interests of our target audience.
Let us denote the user query as $\mathcal{Q}$, the dataset to be analyzed as $\mathcal{D}$, the node set whose interrelationship is of interest as $\mathcal{X}$, the treatment variable as $\mathcal{A}$, the response variable as $\mathcal{Y}$, the mediator variables as $\mathcal{M}$, and the condition of the subgroup of interest as $\mathcal{S}$. The five causal tasks are listed as follows and summarized in Table 1.
@@ 3 Problem Formulation
Task Dataset Nodes Treatment Response Mediator Condition Function Output Format
($\mathcal{D}$) ($\mathcal{X}$) ($\mathcal{A}$) ($\mathcal{Y}$) ($\mathcal{M}$) ($\mathcal{S}$)
CGL Causal Graph ($\mathcal{G}$)
ATE Numeric Value
HTE Numeric Value
MA Numeric Values
OPO Treatment Level
@@
[CEL] Mediation Effect Analysis (MA): Mediators, which act as channels for transmitting treatment effects to the response, are prevalent in various real-world applications. For example, an individual’s mood can mediate the relationship between their work environment and physical health, whereas a negative work environment increases stress, leading to a decline in physical health. In order to facilitate better decision-making, it is critical to address the mediators correctly to obtain accurate effect estimates and reveal the complete causal pathway from treatment to outcome. With the provided dataset, and given the $\mathcal{A}$, $\mathcal{Y}$, and $\mathcal{M}$ specified in $\mathcal{Q}$, our objective is to apply appropriate MA estimators to calculate both the direct effect $\mathbb{E}[\mathcal{Y}(1,\mathcal{M}(0))-\mathcal{Y}(0)]$ and the indirect effect $\mathbb{E}[\mathcal{Y}(1)-\mathcal{Y}(1,\mathcal{M}(0))]$. Here, $\mathcal{M}(\cdot)$ is analogously defined to $\mathcal{Y}(\cdot)$, and $\mathcal{Y}(a,\mathcal{M}(a^{\prime}))$ denote the response that would be observed if treatment $a$ was applied and the mediator set to the level that would be observed if treatment $a^{\prime}$ was applied.
@@
To tackle the aforementioned diverse tasks using a single LLM, we introduce LLM4Causal. By fine-tuning pre-trained LLMs, LLM4Causal is capable of comprehending causal queries, applying appropriate causal tools to analyze the provided tabular dataset, and providing answers by interpreting numerical results in straightforward and fluent language. More technical details of LLM4Causal is discussed in the next section.
4 Proposed Method
In this paper, we introduce a novel three-stage framework to address causality-related tasks, as illustrated in Figure  1. The initial stage involves employing the fine-tuned large language model, LLM4Causal, to interpret users’ inquiries of causal questions by projecting the inquiries into structured templates suitable for functional calls. Specifically, LLM4Causal is carefully fine-tuned with our proposed benchmark dataset, Causal-Retrieval-Bench, enabling it to achieve better performance than other LLMs in classifying causality-related tasks and retrieving relevant information, such as variable names and values. In the second stage, the framework leverages existing tools based on the structured templates generated in the first stage to produce quantitative results for specified causality tasks. However, the numerical outcomes produced in the second stage might not be easily interpretable to a general audience that is unfamiliar with causal inference concepts. To bridge this gap, the third stage of our framework utilizes LLM4Causal to interpret the outputs. To mitigate common pitfalls of LLMs in interpreting numerical outputs, including hallucination and information omissions, the LLM4Causal is further fine-tuned on our proposed Causal-Interpret-Bench dataset. By translating the results from functional calls into clear and easily understandable natural language interpretations, the LLM4Causal addresses the users’ original input queries and provides the interpretation as the final output of the framework.
4.1 STEP1. Interpret User Request
The purpose of the first step is to translate user questions into a causal task classification and a list of related information through two integrated substeps: i) problem classification and ii) attribute extraction. Initially, problem classification categorizes the input question into one of the causal tasks (CGL, ATE, HTE, MA, or OPO) by inferring the underlying user intention. Following this, attribute extraction, conditioned on the task classification, extracts attribute values that are necessary for the expert tools to address the causal problem. Instead of formulating it as a sequential process, we propose to merge them into one sequence-to-sequence (seq2seq) procedure where the output is a structured JSON with the "causal_problem" key and other task-specific keys, e.g. "dataset" ($\mathcal{D}$), "treatment" ($\mathcal{A}$), "outcome" ($\mathcal{Y}$), as keys for the corresponding task. To keep the sequential nature of the problem, we always keep the "causal_problem" key in the first place. The required keys for each causal task are learned implicitly by the parameters of LLM. Some simulated input queries with the corresponding JSON output are shown in Table 2.
@@ 4 Proposed Method
Using the first pair of examples shown in Table 2 as an illustration, the output JSON has three different pre-set keys conditioned on its related causal task, i.e., Causal Graph Learning (CGL), which is the first object to be inferred.Meanwhile, explicit references to key values are required for the remaining keys, such as "disaster_risk_reduction.csv" for the "dataset" key and [’building_code_compliance_rate’, ’disaster_preparedness_campaigns’] for the "nodes" key.
Even with careful prompt engineering, it is challenging to directly adopt a pre-trained LLM for such a seq2seq task, where the latest GPT4 checkpoint still has a 23% error rate (see Section 5.2.1 for details). Thus, considering the deployment cost and the future potential of adding more features, we fine-tuned a local LLM with an augmented dataset consisting both input queries and output JSONs.
4.1.1 Construct Causal-Retrieval-Bench
To collect LLM-generated data, the common approach starts from writing several demonstrated examples. Those demonstrations are then fed to an LLM (e.g., GPT-4) for paraphrasing. We have observed that such an approach suffers from (1) non-controllable topics and variables of interest, which results in content homogeneity, (2) incomplete/non-compliant JSON outputs, and (3) inappropriate paraphrase, where the causal question is transformed to be association-related. To resolve the pain points and prevent data contamination, we propose a three-course procedure, which is illustrated in Figure 3.
@@ 4.1.1 Construct Causal-Retrieval-Bench
Figure 3: Causal-Retrieval-Bench construction procedures for the first step. GPT prompts used in this section can be found in Appendix A. In the left panel, GPT is prompted to list out different topics and measurable variables for each topic. With randomly drawn tasks from section 3 and some variables under the same topic, we generated JSON output in the middle panel. The colored boxes in the left and middle panels are shared topics and variable names of interests. In the right panel, the numbered boxes in template 2 mean blank places to be filled with the task description, the supplied demonstration, JSON inputs, and output restrictions. Common connectors and general instruction in the prompt 2 are shortened into "xxxxx" due to the figure limit.
The procedure can be summarized as an output-first strategy, where the major goal is to improve the correctness and the variety of the output JSON first. An instruction prompt is incorporated to construct causal-related questions in different sentence structures. Specifically, we conducted value generation to prepare a topic database, JSON generation by sampling, and question paraphrasing with prompting.
  • Value Generation: the goal of the first course is to get a candidate pool of the interested topic and variables similar to the RapidAPI [^2] https://rapidapi.com/hub hierarchy tree. We prompt the GPT model to generate some potential topics and related variable names. For each variable, we perform a zero-shot classification by GPT to determine the variable class, i.e., whether it is discrete or continuous. All the collected information is maintained in a three-level hierarchy structure (topic -> variable -> class).
  • JSON Generation: Based on the assumption that causality tasks often involve variables within the same topic, we randomly generate structured JSON data following the topic-variable-class structure. For example, for an average treatment effect estimation task, a topic name is firstly drawn from the stem of the tree as the data file name, and two variables underneath are then considered as candidates for the synthetic treatment and outcome. For causal tasks where a specific group variable is needed, we sampled a random value, either continuous or discrete, based on its class.
  • Question Generation: For each JSON output, we prompt the GPT model to generate five related questions given the context. Utilizing a univariate prompt template, we transform each JSON into an information-augmented prompt consisting of task-specific descriptions, expert-written demo examples, JSON information, and paraphrasing restrictions. Specifically, we implement GPT prompts in this step by a modulized template to improve components’ reusability. Structurally speaking, the prompt starts with a description of the question generation task and the corresponding causal task. It is then followed by information extracted from a JSON from the previous course as the contextual component for the generated question. Subsequently, randomly sampled demonstrations are included as the sample illustrations and output restrictions are itemized as the final remark.
During development, the aforementioned three courses are executed sequentially with quality control on the correctness of generated entities/queries in between. Confused with other statistical concepts, such as correlation and association, the current ChatGPT model has suboptimal performance on the causality-related task. Hence, domain experts are involved after each step to improve the data quality with human annotation. As a result, for the Causal-Retrieval-Bench, we collect 1,500 pairs, i.e., 300 for each causal task, of causal questions with the corresponding JSON outputs.
4.1.2 Fine-tuning a Large Language Model
$\ell(\mathbb{G},\mathcal{S}^{\prime})$ $=$ $\frac{1}{S}\sum_{i=1}^{S}\ell(\mathbb{G}(q_{i}),j_{i})$
$=$ $\frac{1}{SL}\sum_{i=1}^{S}\sum_{k=1}^{L}\textrm{cross-entropy}(\mathbb{G}(q_{i})_{k},j_{ik})$
$=$ $\frac{1}{SL}\sum_{i=1}^{S}\sum_{k=1}^{L}\sum_{l=1}^{V}-I(j_{ik}=w_{l})log(P(\mathbb{G}(q_{i})_{k})=w_{l}),$
4.2 STEP2. Assign and Execute Causal Tools
In this stage, to address user queries, we first select suitable causal learning algorithms based on the task class determined in Step 1, as depicted in Figure 4. Following this selection, the algorithm is executed automatically, using both the extracted information from Step 1 and the user-provided dataset as inputs, to get the estimated result. Depending on the causal task, the output result can be a causal graph $\mathcal{G}$, numerical values of estimated causal effect (estimated expected value of $\mathcal{Y}$), or a treatment level $a\in\mathcal{A}$. LLM4Causal facilitates a variety of approaches for each causal task by leveraging well-known public causal packages, including econml(Oprescu et al., [n. d.]), causalml(Chen et al., 2020), causal-learn(Zheng et al., 2023), and CausalDM(Song et al., 2023). Additionally, we remark that LLM4Causal is designed with flexibility, allowing for the incorporation of new methodologies through the simple integration of function scripts. In the current version, LLM4Causal randomly selects a method for analysis. However, we are in the process of enhancing LLM4Causal for a subsequent release, which aims to provide i) a more interactive platform where users can specify their preferred methods; ii) a comprehensive causal analysis report, offering outputs from all methods with detailed comparisons upon request; and iii) an advanced method selection feature for automatically determining the most suitable approach.
Figure 4: Causal Tool Assignment and Execution in the Second Step
4.3 STEP3. Interpret the Numerical Output
As is discussed in Section 4.2, the function call in our pipeline produces a numerical value of causal effect, a recommended action level, or an adjacency matrix indicating causal relations in a causal graph. Interpreting these outputs needs domain expertise and may not be easily understood by a general audience lacking a causal background. To enhance user-friendliness and reduce the learning curve of LLM4Causal, our final stage entails the interpretation of the function output using fluent, natural language. The process starts with a template interpretation that automatically converts the numerical output to a one-sentence summary (See Appendix B.1 for examples). The final human-readable interpretation is then produced through the utilization of a calibrated LLM, which uses the template summary and other information obtained from previous steps as input and is guided by a designed prompt. The prompt instructs the LLM to avoid generating hallucinatory content, ensure the inclusion of all relevant information, and provide cohesive responses in the relevant context of the original question (See Appendix B.2 for details).
Admittedly, we may directly take the crafted prompt and let a pre-trained LLM such as GPT-4 generate interpretation for us. However, the interpretations obtained this way are problematic. Human annotations indicate that about 25% of the outputs exhibit hallucinations, 10% contain missing information, and 35% lack fluency (see their definition below ). Therefore, we calibrated the same LLM model with additional human annotated interpretation for more accurate result interpretation and the prompt includes the aforementioned entities as basic elements and a graphical illustration can be found in figure 5.
Below, we outline the process used to generate the golden dataset Causal-Interpret-Bench. The approach for model fine-tuning is similar to that used in step 1 and is omitted here for brevity.
Figure 5: Illustration of building Causal-Interpret-Bench in the Third Step. Details about used prompts can be found in Appendix B. In template 3 we fixed the introduction and the final restriction parts because the instruction is about interpretation, and the remaining contextual information is filled differently every time when calling the GPT API. Similar to the idea of section 4.1.1, the general introduction and the restriction on outputs are reused across different causal problems. We feed the original causal query, the causal task associated with the query, and the employed methodology with corresponding function outputs to the prompt as the interpretation context.
4.3.1 Build Causal-Interpret-Bench
Likewise to the procedures in Step 1, we first use the GPT model to generate a silver dataset of interpretations, employing the templates in Appendix B.1 and prompts in Appendix B.2. Among the required inputs, user queries and their corresponding causal task classifications from section 4.1.1 are reused. Using the methodology described in 4.2, we paired the existing information with a causal function. As for the function outputs, the numerical values are randomly generated in formats specific to each task. To create the final golden dataset, we then have annotators manually revise 400 interpretations, adhering to the evaluation rubric outlined below:
  • Incompleteness: The interpretation lacks one or more essential components, such as the dataset, method, results, and/or all involved variables.
While there are diverse metrics developed to evaluate the factual consistency between the output summaries/interpretation and the original context Zhang et al. (2019) or to assess the conciseness and readability of output summaries, such as the ROUGE score Lin (2004), they fall short of our specific needs. The interpretation we require is highly specialized, necessitating a clear distinction between causality and association, guaranteeing that no crucial information is missed, and ensuring it is comprehensible to general audiences, among other requirements. Current automatic evaluation based on sentence similarity metrics fail to adequately incorporate human preferences, making them unsuitable for our application. Therefore, we carefully defined the three aforementioned rubrics instead to guide the human annotation.
5 Experiments
5.1 End-to-End Performance Evaluation: A Case Study
The user experience of our work is demonstrated in figure 6, 7, and 8, which presents the capabilities of answering causal-related questions with LLM4Causal as a chatbot. For each of the CGL, ATE, and HTE tasks, we show one end-to-end example. The demo website is developed by the Streamlit [^3] https://streamlit.io/ package in Python with the intention that users can chat with our model as a ChatGPT-like environment with causal problem-solving ability. All three examples use a publicly available dataset called "cancer_1000_1.csv" from Li and Beek (2018).
Figure 6: Demo Example of Answering Causal Structure Learning Question
Figure 6 shows the system results (including intermediate steps) for the question "What is the causal relationship between smoking and cancer in the cancer_1000_1.csv dataset?" and LLM4Causal successfully identified it as a CGL task with the JSON output "’causal_problem’: [’CSL’, ’CGL’], ’dataset’: [’cancer_1000_1.csv’], ’covariates’: [None]". The screenshot does not capture them due to the web page limit. After the JSON step (step 1 in 4.1), the service loaded the data from the hard drive and triggered a PC algorithm to discover the causal relationship between the interested variables. Finally, the original questions were answered in natural language, using the results from the graph_learning function.
Figure 7: Demo Example of Answering Average Treatment Effect Question
Figure 8: Demo Example of Answering Heterogeneous Treatment Effect Question
Both figure 7 and figure 8 showcase the full conversation between the user and LLM4Causal. The difference between these two examples is if we wanted the effect estimation limited to a specific group of populations. If so (for example pollution = "high") then we need to trigger the heterogeneous treatment effect instead of the average treatment effect with an additional input "condition". We can see that LLM4Causal discriminated different problems prosperously and provided reasonable interpretations based on the provided context.
5.2 Ablation Analysis
We also carefully evaluated the model performance with an ablation analysis to determine the effectiveness of the causal entity extraction described in section 4.1. Our fine-tuned model is benchmarked with a GPT4 variant (with or without function call features). As a further step, more simulation studies involving end-to-end and interpretation evaluation will be released soon.
As for LLM4Causal, we trained two different checkpoints: ‘LLM4Causal-Retrieve’ only trains on Causal-Retrieval-Benchand ‘LLM4Causal-Mixed’ or ‘LLM4Causal’ simultaneously trained on both datasets. The purpose of showcasing ‘LLM4Causal-Mixed’ and ‘LLM4Causal-Retrieve’ is to show that mixing both sources of the augmented data will not harm the inference quality and that only having one univariate model is sufficient.
5.2.1 Causal Entity Extraction
We left 150 samples of the training data from the process of section 4.1. For each simulated causal question, the benchmarked model was asked to produce a JSON output following the structure rule for the downstream causal function, and the output was validated by the ground truth label by comparing the value for each key. The accuracy scores for correctly specifying the causal task classification and exactly matching the values under each key are reported. As a consequence of different causal tasks having variant keys, if the model misclassifies the causal task in the first place, its performance for the corresponding keys to the problem will be harmed.
Here we are benchmarking GPT4 with the function call feature enabled. It was developed by OpenAI in 2023 and demonstrated a strong capability of zero-shot function selection and argument extraction. Multiple papers (Qin et al., 2023b; Zhang, 2023) treated its output as the golden label for local LLM development. We created one description file for each causal task with the specification of required input arguments, the output was mapped into one JSON by with an additional "causal_problem" key storing the selected function.
Table 3: Causal Entity Extraction Performance
Causal Task Dataset Nodes Treatment Response Mediator Condition
gpt3.5-turbo 0.77 0.77 0.27 0.60 0.61 0.81 0.33
gpt3.5-turbo (guided) - 1.00 0.27 0.84 0.80 0.94 0.42
gpt4-turbo 0.77 0.77 0.27 0.67 0.62 0.67 0.32
gpt4-turbo (guided) - 1.00 0.27 0.90 0.86 0.97 0.40
LLM4Causal-Retrieve 0.98 0.95 0.81 0.97 0.98 0.94 0.97
LLM4Causal-Mixed 0.99 0.99 0.88 0.98 0.97 1.00 0.98
It is shown in the table 3 that the turbo version of both GPT3.5 and GPT4 can perform the causal task identification task relatively well, but their accuracy on detailed argument extraction is limited, especially for keys with complicated formats such as condition and nodes. As a further step, we only provided the correct function to the model (marked as guided) and tried to investigate if the performance was masked by selecting the incorrect causal task. We can see significant improvement as expected. For gpt4 specifically, we had an increase in the dataset ($0.77\rightarrow 1.00$), treatment ($0.67\rightarrow 0.90$), response ($0.62\rightarrow 0.86$), and mediator ($0.67\rightarrow 0.97$), but it still made considerable mistakes on exactly matching nodes and conditions. As for LLM4Causal, we can see that both checkpoints can solve such a seq2seq problem very well, only the Condition key having an exact match accuracy rate that is less than $0.9$. A seemingly surprising fact is the mixed version even outperforms the retrieve-only version on most tasks except the nodes, which may be due to the inclusion of the interpretation data, which also involves causal queries and their corresponding causal tasks with interpretations, increasing the word connection between causal tasks and their meanings.
6 Conclusion
In this paper, we have proposed LLM4Causal, the first end-to-end user-friendly large language model with causal decision-making ability. Without requiring in-depth domain knowledge, the proposed model could be easily used for the general audience, which remedies the weakness of the current LLM application on the causal tasks. Furthermore, we have proposed a novel data generation pipeline by utilizing GPT-based LLM models and human annotations to improve the data quality in terms of both data variety and accuracy. Fine-tuned to interpret users’ query and function outputs, LLM4Causal has shown superior performance in causal entity extraction. As a further step, more simulation studies involving end-to-end and interpretation evaluation will be released soon.
To conclude, we briefly discuss some limitations followed by inspired possible future directions. First, the proposed model is only capable of handling five major causal tasks, but it is worth mentioning that our framework could be easily extended to support more causal tasks with cutting-edge methodologies. Second, the current approach randomly selects one causal tool in step 2, but it could be a valuable direction to enable the model to automatically select the best available tools based on data characteristics and context information. Third, in this project, we exploit the LLM’s capability of using statistical tools to conduct data analysis, while the internal knowledge that LLM obtains during its training may also be beneficial. It could be of great value to integrate LLM’s tool-usage ability with LLM’s internal knowledge to further improve its performance on causal tasks. Lastly, a potential direction to explore is the interactive potential of the framework so that the proposed method could iteratively adapt to users’ feedback, which could be beneficial in following the users’ preferences for causal tools, encouraging users to collect necessary information when their input is not enough, and enhancing the interpretations based on their extra input in the problem context.
Appendix A Details about Step 1
@@
Given the details below, generate five diverse and plain-worded questions [Task requirements]:
@@
Given the details below, generate five diverse and plain-worded questions asking about the existence
@@
- Is there a method to discover every direct influence present in the employee_data.csv dataset?
- Within the employee_data.csv dataset, how many instances of one factor directly causing another
@@
- How many causal connections can be identified in the employee_data.csv dataset?
@@
- Interested variable: education (edu), job satisfaction (job_satisfaction)
@@
- Within the job_data.csv dataset, is there causal links between education (edu) and job satisfaction
@@
- Does the job_data.csv data reveal any direct relationships between education (edu) and job
@@
- Are there discernible causal links between education (edu) and job satisfaction (job_satisfaction)
in the job_data.csv dataset?
@@
- Interested variable: temperature (temp), humidity (humid), rainfall (rain), local latitude (lat),
local altitude (alt)
@@
- What are causal links among temperature (temp), humidity (humid), rainfall (rain), local latitude
(lat), and local altitude (alt) in dataset weather_data.csv?
- Within the dataset weather_data.csv, what is the nature of the causal links among temperature (temp),
humidity (humid), rainfall (rain), local latitude (lat), and local altitude (alt)?
- In the weather_data.csv dataset, what causal connections can be identified among temperature (temp),
humidity (humid), rainfall (rain), local latitude (lat), and local altitude (alt)?
Ensure that the questions employ a mix of different phrasing and diverse sentence structures.
@@
Ensure that the questions place a strong emphasis on the effect existence or the total number.
Ensure that all variable names under interested variables field exist in all five quetsion.
Ensure that the provided interested variable names are integrated naturally into the questions without
discarding or altering any part of them.
Appendix B Details about Step 3
B.1 Template Outcome

### 12 Apr 2024 v3

@@ A.4 Large Language Model with Tool Usage
Appendix B Detailed Problem Formulation
In this paper, we aim to augment existing pre-trained LLMs with the proficiency to address causal inquiries, thereby opening the door to causal decision-making processes for general audiences. As introduced previously, causal decision-making tasks can be classified into three categories: 1) causal structure learning (CSL), 2) causal effect learning (CEL), and 3) causal policy learning (CPL). Each of these primary categories consists of multiple tasks, and our study focuses on five key tasks that are particularly relevant to the interests of our target audience. Let us denote the user query as $\mathcal{Q}$, the dataset to be analyzed as $\mathcal{D}$, the node set whose interrelationship is of interest as $\mathcal{X}$, the treatment variable as $\mathcal{A}$, the response variable as $\mathcal{Y}$, the mediator variables as $\mathcal{M}$, and the condition of the subgroup of interest as $\mathcal{S}$. The five causal tasks are listed as follows and summarized in Table 5.
@@ Appendix B Detailed Problem Formulation
Task Dataset Nodes Treatment Response Mediator Condition Function Output Format ($\mathcal{D}$) ($\mathcal{X}$) ($\mathcal{A}$) ($\mathcal{Y}$) ($\mathcal{M}$) ($\mathcal{S}$) CGL Causal Graph ($\mathcal{G}$) ATE Numeric Value HTE Numeric Value MA Numeric Values OPO Treatment Level
@@
[CEL] Mediation Effect Analysis (MA): Mediators, which act as channels for transmitting treatment effects to the response, are prevalent in various real-world applications. With the provided dataset, and given the $\mathcal{A}$, $\mathcal{Y}$, and $\mathcal{M}$ specified in $\mathcal{Q}$, MA’s objective is to calculate both the direct effect $\mathbb{E}[\mathcal{Y}(1,\mathcal{M}(0))-\mathcal{Y}(0)]$ and the indirect effect $\mathbb{E}[\mathcal{Y}(1)-\mathcal{Y}(1,\mathcal{M}(0))]$. Here, $\mathcal{M}(\cdot)$ is analogously defined to $\mathcal{Y}(\cdot)$, and $\mathcal{Y}(a,\mathcal{M}(a^{\prime}))$ denote the response that would be observed if treatment $a$ was applied and the mediator set to the level that would be observed if treatment $a^{\prime}$ was applied.
@@
To tackle the aforementioned diverse tasks using a single LLM, we introduce LLM4Causal. By fine-tuning pre-trained LLMs, LLM4Causal is capable of comprehending causal queries, applying appropriate causal tools to analyze the provided tabular dataset, and providing answers by interpreting numerical results in straightforward and fluent language. More technical details of LLM4Causal is discussed in Section 3.
Appendix C Supplimentary Methodology
C.1 Demonstration input-output pairs
Following procedures described in 3.1, which condition on the task classification, extract attribute values that are necessary for the expert tools to address the causal problem. For each input question, the corresponding output is a structured JSON with the "causal_problem" key and other task-specific keys, e.g. "dataset" ($\mathcal{D}$), "treatment" ($\mathcal{A}$), "outcome" ($\mathcal{Y}$). Some simulated input queries with the corresponding JSON output are shown in Table 6.
@@ C.1 Demonstration input-output pairs
C.2 Details about Step 1
Figure 4 is a detailed version of figure 3, wherein the caption we describes how such a process goes.
@@ C.2 Details about Step 1
Figure 4: Causal-Retrieval-Bench construction procedures for the first step. GPT prompts used in this section can be found in Appendix C.2. In the left panel, GPT is prompted to list out different topics and measurable variables for each topic. With randomly drawn tasks from section 2 and some variables under the same topic, we generated JSON output in the middle panel. The colored boxes in the left and middle panels are shared topics and variable names of interests. In the right panel, the numbered boxes in template 2 mean blank places to be filled with the task description, the supplied demonstration, JSON inputs, and output restrictions. Common connectors and general instruction in prompt 2 are shortened into "xxxxx" due to the figure limit.
@@
Given the details below, generate five diverse and plain-worded questions
[Task requirements]:
@@
Given the details below, generate five diverse and plain-worded questions
asking about the existence
@@
- Is there a method to discover every direct influence present in
the employee_data.csv dataset?
- Within the employee_data.csv dataset, how many instances of one factor
directly causing another
@@
- How many causal connections can be identified in the
employee_data.csv dataset?
@@
- Interested variable: education (edu), job satisfaction
(job_satisfaction)
@@
- Within the job_data.csv dataset, is there causal links between education
(edu) and job satisfaction
@@
- Does the job_data.csv data reveal any direct relationships between
education (edu) and job
@@
- Are there discernible causal links between education (edu) and job
satisfaction (job_satisfaction) in the job_data.csv dataset?
@@
- Interested variable: temperature (temp), humidity (humid), rainfall
(rain), local latitude (lat), local altitude (alt)
@@
- What are causal links among temperature (temp), humidity (humid),
rainfall (rain), local latitude (lat), and local altitude (alt) in dataset
weather_data.csv?
- Within the dataset weather_data.csv, what is the nature of the causal
links among temperature (temp), humidity (humid), rainfall (rain),
local latitude (lat), and local altitude (alt)?
- In the weather_data.csv dataset, what causal connections can be
identified among temperature (temp), humidity (humid), rainfall (rain),
local latitude (lat), and local altitude (alt)?
Ensure that the questions employ a mix of different phrasing and diverse
sentence structures.
@@
Ensure that the questions place a strong emphasis on the effect existence
or the total number.
Ensure that all variable names under interested variables field exist in
all five quetsion.
Ensure that the provided interested variable names are integrated
naturally into the questions without discarding or altering any part of them.
Specifically, we implement GPT prompts in this step by a modulized template to improve components’ reusability. Structurally speaking, the prompt starts with a description of the question generation task and the corresponding causal task. It is then followed by information extracted from a JSON from the previous course as the contextual component for the generated question. Subsequently, randomly sampled demonstrations are included as the sample illustrations and output restrictions are itemized as the final remark.
Using the example of CGL task in Table 6 as an illustration, the output JSON has two task-specific keys conditioned on its inferred classification as a CGL task: "dataset" with the value "disaster_risk_reduction.csv," and "nodes", which includes ’building_code_compliance_rate’ and ’disaster_preparedness_campaigns’.
C.3 Details about Step 2
Figure 5 describes the procedure of step 2 workflow, where we utilize the output from step 1 to select an appropriate input function as well as its necessary input fields to conduct causal analysis.
Figure 5: Causal Tool Assignment and Execution in the Second Step. The algorithm is executed automatically, using both the extracted information from Step 1 and the user-provided dataset as inputs, to get the estimated result
C.4 Details about Step 3
C.4.1 Template Outcome